the backyard of a neighbor's house is rectangular, and its length is 5 feet more than its width. if the area of the yard is 500 square feet, how much fencing would be required to enclose the yard
Answer by TimothyLamb(4379)

---
x = y + 5
xy = 500
---
xy = 500
(y + 5)y = 500
yy + 5y - 500 = 0
---
the above quadratic equation is in standard form, with a=1, b=5, and c=-500
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 5 -500
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
y = 20
y = -25
---
negative width doesn't make sense so use the positive root:
---
answer:
w = 20 ft
L = 25 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php